Integrativelong dashOptimal capital structure   The board of directors of Morales​ Publishing, Inc., has commissioned a capital structure study. The company has total assets of $40,800,000 . It has earnings before interest and taxes of $7,910,000 and is taxed at a rate of 27% .

a. Create a spreadsheet showing values of debt and equity as well as the total number of​ shares, assuming a book value of  $25 per share.

b. Given the​ before-tax cost of debt at various levels of​ indebtedness, calculate the yearly interest expenses.

c. Using EBIT of $7,910,000 ​, a 27% tax​ rate, and the information developed in parts ​(a​) and ​(b​)​, calculate the most likely earnings per share for the firm at various levels of indebtedness. Mark the level of indebtedness that maximizes EPS.

d. Using the EPS developed in part ​(c​)​, the estimates of required​ return, r Subscript s and the equation Upper P0=EPS/ rs ​, estimate the value per share at various levels of indebtedness. Mark the level of indebtedness in the following table that results in the maximum price per​ share, P0 .

e. Prepare a recommendation to the board of directors of Morales Publishing that specifies the degree of indebtedness that will accomplish the​ firm's goal of optimizing shareholder wealth. Use your findings in parts ​(a​) through ​(d​) to justify your recommendation. a. Create a spreadsheet showing values of debt and equity as well as the total number of​ shares, assuming a book value of  $ 25 per share.
Answer #1
Solution
a) Book Value/Share $25
A B C D E
C=A*B D=B-C E=D/25
Debt Total Assets Debt Equity No of Shares
0% 40800000 0 40800000 1632000
10% 40800000 4080000 36720000 1468800
20% 40800000 8160000 32640000 1305600
30% 40800000 12240000 28560000 1142400
40% 40800000 16320000 24480000 979200
50% 40800000 20400000 20400000 816000
60% 40800000 24480000 16320000 652800
70% 40800000 28560000 12240000 489600
80% 40800000 32640000 8160000 326400
90% 40800000 36720000 4080000 163200
b) Interest on debt is assumed as 10%
EBIT=7910000
A B C D
C=A*B D=C*10%
Debt Total Assets Debt Interest Expense
0% 40800000 0 0
10% 40800000 4080000 408000
20% 40800000 8160000 816000
30% 40800000 12240000 1224000
40% 40800000 16320000 1632000
50% 40800000 20400000 2040000
60% 40800000 24480000 2448000
70% 40800000 28560000 2856000
80% 40800000 32640000 3264000
90% 40800000 36720000 3672000
c) EBIT=7910000
A B C D E F G H I
C=A*B D=C*10% E=EBIT-D =E*27% G=E-F I=G/H
Debt Total Assets Debt Interest Expense EBT Tax@27% EAT No of shares EPS
0% 40800000 0 0 7910000 2135700 5774300 1632000 3.54
10% 40800000 4080000 408000 3830000 1034100 2795900 1468800 3.73
20% 40800000 8160000 816000 -250000 -67500 -182500 1305600 3.97
30% 40800000 12240000 1224000 -4330000 -1169100 -3160900 1142400 4.27
40% 40800000 16320000 1632000 -8410000 -2270700 -6139300 979200 4.68
50% 40800000 20400000 2040000 -12490000 -3372300 -9117700 816000 5.25
60% 40800000 24480000 2448000 -16570000 -4473900 -1.2E+07 652800 6.11
70% 40800000 28560000 2856000 -20650000 -5575500 -1.5E+07 652800 5.65
80% 40800000 32640000 3264000 -24730000 -6677100 -1.8E+07 652800 5.20
90% 40800000 36720000 3672000 -28810000 -7778700 -2.1E+07 652800 4.74
EBIT =Earnings Before Interest and Tax
EBT=Earnings After Interest Before Tax
EAT=Earnings After Tax
d) Estimated Required return assumed as 15%
A B C D E F
F=E/15%
Debt Total Assets Debt No of shares EPS Po
0% 40800000 0 1632000 3.54 35.38
10% 40800000 4080000 1468800 3.73 37.29
20% 40800000 8160000 1305600 3.97 39.66
30% 40800000 12240000 1142400 4.27 42.72
40% 40800000 16320000 979200 4.68 46.80
50% 40800000 20400000 816000 5.25 52.51
60% 40800000 24480000 652800 6.11 61.08
70% 40800000 28560000 652800 5.65 56.52
80% 40800000 32640000 652800 5.20 51.95
90% 40800000 36720000 652800 4.74 47.39